# Geometric representations of braid and Yang-Baxter gates

### v1

@@ IV.1 Yang-Baxterization
$(R(x)\otimes 1\!\!1_{2})(1\!\!1_{2}\otimes R(xy))(R(y)\otimes 1\!\!1_{2})\\ =(1\!\!1_{2}\otimes R(y))(R(xy)\otimes 1\!\!1_{2})(1\!\!1_{2}\otimes R(x)).$ (20)
It is called the Yang-Baxter equation, where $x$ is called the spectral parameter [71, 72]. The Yang-Baxter equation lays the foundation of the quantum inverse scattering method which systematically solves the Yang-Baxter integrable models [11].
@@
The $R$ matrix obtained from the Yang-Baxterization on the unitary braid matrix with two distinct eigenvalues is unitary if the spectral parameter is real and $B^{2}\neq 1\!\!1$.
The proof is straightforward and can be found in Appendix A. If $B^{2}=1\!\!1$, such as $B=\text{SWAP}$, the Yang-Baxterization gives a trivial $R(x)$ matrix where the spectral parameter $x$ is only shown in the global phase.
IV.2 Yang-Baxter gates from Yang-Baxterization
@@
It suggests that they have the same eigenvalues, denoted as
@@
$R_{l,3}(x)=$ $\alpha(x,\lambda_{l,2})B_{I}+\gamma(x,\lambda_{l,1})B^{\dagger}_{l},$ (34b)
@@
Specifically, $R_{I,1}$ is the $R$ matrix of the XXZ model [77]. The braid gate parameter in $R_{I,1}$ corresponds to the anisotropic coupling of the spin chain. The $R_{III,1}$ matrix can also be obtained from the trigonometric limits of elliptic $R$ matrix [78]. Moreover, the $R_{III,1}$ matrix is locally equivalent to the $R$ matrix of the sine-Gordon model [79]. The matrix expressions of $R_{l,1}$ $R_{l,2}$, and $R_{l,3}$, namely total nine $R$ gates, can be found in Appendix B.
@@
The optimal gate and pulse constructions of $R_{IV}(\chi)$ with $\varphi_{1}=0$ have been detailed studied in [64].
IV.3 Geometric representations of Yang-Baxter gates
@@
The Yang-Baxter gates $R_{I,2}$ and $R_{I,3}$ are SWAP-like gates (exchanging states $|01\rangle$ and $|10\rangle$ with some phases). See the Appendix B. Therefore, they have similar geometric representations as $B_{I}$. Specifically, they have the nonlocal parameters
@@
Compare with the nonlocal parameters of $R_{I,1}$ in Eq. (37), then we find that $R_{I,1}$ and $R_{III,1}$ have the correspondence through $2\varphi_{1}\leftrightarrow\varphi$ and $2\mu\leftrightarrow\mu$. Therefore, $R_{III,1}$ is also on the faces $OA_{1}A_{2}$ and $OA_{2}A_{3}$ of the two-qubit tetrahedron. See Fig. 4.
@@
IV.4 Gate decompositions of Yang-Baxter gates
@@
Based on the nonlocal parameters of the Yang-Baxter gates, we find the optimal gate decomposition (with the minimal number of CNOTs) of $R_{I}$ and $R_{III}$ in terms of the universal gate set $\{R_{z}(\theta),H,\text{CNOT}\}$. See Figs. 5 and 6. For $R_{I}$, we define the parameter $\omega$ as
@@
which only appears in the single-qubit gates of the decomposition. It is also called the $q$ deformation parameter of the $R$ matrix [77].
@@
The analysis on $R_{II}$ to be the Clifford gate is the same as $R_{I}$ due to the relation (49). For the case $R_{III}$, we see that there are $T$ gates in the decompositions. See Fig. 6. Therefore, the necessary and sufficient condition for $R_{III}$ to be the Clifford gate becomes $\varphi_{2}\in\{k\pi+\pi/2,k\in\mathbb{Z}\}$, and $a_{III,1},c_{III,1},\varphi_{III,2},\varphi_{III,3}\in\{k\pi/2,k\in\mathbb{Z}\}$. Specifically, the condition $a_{III,1},c_{III,1}\in\{k\pi/2,k\in\mathbb{Z}\}$ (for $R_{III,1}$ to be the Clifford gate) is satisfied if $\mu=0$ or $\varphi_{1}\in\{k\pi/2,k\in\mathbb{Z}\}$. It is similar with $R_{I,1}$. When $\mu=0$ or $\varphi_{1}\in\{k\pi+\pi/2,k\in\mathbb{Z}\}$, we have the Clifford $R_{III,2}$, which is locally equivalent to iSWAP gate; when $\varphi_{1}\in\{k\pi,k\in\mathbb{Z}\}$, we have the Clifford $R_{III,2}$, which is locally equivalent to SWAP gate. Similar results are applied to $R_{III,3}$. The Yang-Baxter gate $R_{IV}$ is Clifford gate if $\varphi_{1}\in\{k\pi,k\in\mathbb{Z}\}$ and $\chi\in\{k\pi/4,k\in\mathbb{Z}\}$, which is locally equivalent to CNOT. We summary the results in Table 2.
@@ IV.5 Entangling powers
The entangling power of the two-qubit gate defined in Eq. (9) quantifies the ability to generate entanglement from the product state. It can be viewed as the interacting strength between qubits in the brick-wall circuits [51]. Given the nonlocal parameters of $R_{I}$, we numerically calculate its entangling powers in terms of the braid gate parameter $\varphi$ and the spectral parameter $\mu$. See Fig. 7. When $\mu=0$ or $\varphi\in\{k\pi,k\in\mathbb{Z}\}$, $R_{I,1,2,3}$ has the zero entangling power, corresponding to the trivial two-qubit gate or SWAP gate. When $\mu\rightarrow\infty$ and $\varphi\in\{k\pi+\pi/2,k\in\mathbb{Z}\}$, $R_{I,1,2,3}$ has the maximal entangling power, corresponding to the iSWAP gate. Note that $\mu\leftrightarrow-\mu$ does not change the magnitude of the nonlocal parameters and the entangling power, therefore we only draw $\mu\geq 0$ in Fig. 7.
@@
V Conclusions
In this work, we have studied the geometric representations of braid and Yang-Baxter gates, where the Yang-Baxter gates are obtained from the Yang-Baxterization of braid gates. Braid and Yang-Baxter gates on the same point of the two-qubit tetrahedron can be transformed from each other through single-qubit gates. In the two-qubit tetrahedron representation, we find that the Yang-Baxter gates can only exist on the edges $A_{2}A_{3}$ and $OA_{1}$, and on the faces $OA_{2}A_{3}$ and $A_{1}A_{2}A_{3}$. Based on their geometric representations, we give their gate decomposition in terms of the universal gate set $\{R_{z}(\theta),H,\text{CNOT}\}$ with the minimal number of CNOTs. We identify the parameters that make the braid and Yang-Baxter gates to be the Clifford gate, the matchgate, and the dual unitary gate. The braid and Yang-Baxter gates have the maximal entangling power if they are locally equivalent to iSWAP or CNOT.
@@ V Conclusions
Acknowledgements.
@@ Appendix A Proof of Theorem 2
$R_{I,1}(\mu)\cong\begin{pmatrix}1&0&0&0\\ 0&\dfrac{\sin\varphi}{\sin(\varphi-i\mu)}&\dfrac{-ie^{i\omega}\sinh\mu}{\sin(\varphi-i\mu)}&0\\ 0&\dfrac{-ie^{-i\omega}\sinh\mu}{\sin(\varphi-i\mu)}&\dfrac{\sin\varphi}{\sin(\varphi-i\mu)}&0\\ 0&0&0&1\\ \end{pmatrix},\qquad R_{II,1}(\mu)\cong\begin{pmatrix}\dfrac{\sin\varphi}{\sin(\varphi-i\mu)}&0&0&\dfrac{-ie^{i\omega}\sinh\mu}{\sin(\varphi-i\mu)}\\ 0&1&0&0\\ 0&0&1&0\\ \dfrac{-ie^{-i\omega}\sinh\mu}{\sin(\varphi-i\mu)}&0&0&\dfrac{\sin\varphi}{\sin(\varphi-i\mu)}\\ \end{pmatrix},$ (59a)
$R_{I,2}(\mu)\cong\frac{1}{\sqrt{\Delta_{I,2}}}\begin{pmatrix}\sinh\left(\frac{1}{2}(\mu+i\varphi)\right)&0&0&0\\ 0&0&e^{i\omega}\sinh\left(\frac{1}{2}(\mu-i\varphi)\right)&0\\ 0&e^{-i\omega}\sinh\left(\frac{1}{2}(\mu-i\varphi)\right)&0&0\\ 0&0&0&\sinh\left(\frac{1}{2}(\mu+i\varphi)\right)\\ \end{pmatrix},$ (59b)
$R_{I,3}(\mu)\cong\frac{1}{\sqrt{\Delta_{I,3}}}\begin{pmatrix}\cosh\left(\frac{1}{2}(\mu+i\varphi)\right)&0&0&0\\ 0&0&e^{i\omega}\cosh\left(\frac{1}{2}(\mu-i\varphi)\right)&0\\ 0&e^{-i\omega}\cosh\left(\frac{1}{2}(\mu-i\varphi)\right)&0&0\\ 0&0&0&\cosh\left(\frac{1}{2}(\mu+i\varphi)\right)\\ \end{pmatrix},$ (59c)
$R_{II,2}(\mu)\cong\frac{1}{\sqrt{\Delta_{II,2}}}\begin{pmatrix}0&0&0&e^{i\omega}\sinh\left(\frac{1}{2}(\mu-i\varphi)\right)\\ 0&\sinh\left(\frac{1}{2}(\mu+i\varphi)\right)&0&0\\ 0&0&\sinh\left(\frac{1}{2}(\mu+i\varphi)\right)&0\\ e^{-i\omega}\sinh\left(\frac{1}{2}(\mu-i\varphi)\right)&0&0&0\\ \end{pmatrix},$ (59d)
$R_{II,3}(\mu)\cong\frac{1}{\sqrt{\Delta_{II,3}}}\begin{pmatrix}0&0&0&e^{i\omega}\cosh\left(\frac{1}{2}(\mu-i\varphi)\right)\\ 0&\cosh\left(\frac{1}{2}(\mu+i\varphi)\right)&0&0\\ 0&0&\cosh\left(\frac{1}{2}(\mu+i\varphi)\right)&0\\ e^{-i\omega}\cosh\left(\frac{1}{2}(\mu-i\varphi)\right)&0&0&0\\ \end{pmatrix},$ (59e)
@@
$R_{III,1}(\mu)\cong\begin{pmatrix}\dfrac{\cosh\mu\cos\varphi_{1}}{\cosh(\mu+i\varphi_{1})}&0&0&-e^{i\varphi_{2}}\dfrac{\sinh\mu\sin\varphi_{1}}{\cosh(\mu+i\varphi_{1})}\\ 0&i\dfrac{\cosh\mu\sin\varphi_{1}}{\sinh(\mu+i\varphi_{1})}&-\dfrac{\sinh\mu\cos\varphi_{1}}{\sinh(\mu+i\varphi_{1})}&0\\ 0&-\dfrac{\sinh\mu\cos\varphi_{1}}{\sinh(\mu+i\varphi_{1})}&i\dfrac{\cosh\mu\sin\varphi_{1}}{\sinh(\mu+i\varphi_{1})}&0\\ e^{-i\varphi_{2}}\dfrac{\sinh\mu\sin\varphi_{1}}{\cosh(\mu+i\varphi_{1})}&0&0&\dfrac{\cosh\mu\cos\varphi_{1}}{\cosh(\mu+i\varphi_{1})}\end{pmatrix},$ (62a)
$R_{III,2}(\mu)\cong\frac{1}{\sqrt{\Delta_{III,2}}}\begin{pmatrix}\sinh\mu\cos\varphi_{1}&0&0&e^{i\varphi_{2}}\cosh\mu\sin\varphi_{1}\\ 0&i\cosh\mu\sin\varphi_{1}&-\sinh\mu\cos\varphi_{1}&0\\ 0&-\sinh\mu\cos\varphi_{1}&i\cosh\mu\sin\varphi_{1}&0\\ -e^{-i\varphi_{2}}\cosh\mu\sin\varphi_{1}&0&0&\sinh\mu\cos\varphi_{1}\\ \end{pmatrix},$ (62b)
$R_{III,3}(\mu)\cong\frac{1}{\sqrt{\Delta_{III,3}}}\begin{pmatrix}\cosh\mu\cos\varphi_{1}&0&0&-e^{i\varphi_{2}}\sinh\mu\sin\varphi_{1}\\ 0&i\sinh\mu\sin\varphi_{1}&-\cosh\mu\cos\varphi_{1}&0\\ 0&-\cosh\mu\cos\varphi_{1}&i\sinh\mu\sin\varphi_{1}&0\\ e^{-i\varphi_{2}}\sinh\mu\sin\varphi_{1}&0&0&\cosh\mu\cos\varphi_{1}\\ \end{pmatrix},$ (62c)

### v2

@@ 4.1 Yang-Baxterization
$(R(x)\otimes 1\!\!1_{2})(1\!\!1_{2}\otimes R(xy))(R(y)\otimes 1\!\!1_{2})=(1\!\!1_{2}\otimes R(y))(R(xy)\otimes 1\!\!1_{2})(1\!\!1_{2}\otimes R(x)).$ (21)
It is called the Yang-Baxter equation, where $x$ is called the spectral parameter [72, 73]. The Yang-Baxter equation is the key in the quantum inverse scattering method which systematically solves the Yang-Baxter integrable models [11].
@@
If the braid gate $B$ has two distinct eigenvalues and satisfying $B^{2}\neq 1\!\!1$, the $R$ matrix obtained from the Yang-Baxterization on $B$ is unitary if the spectral parameter is real.
The proof is straightforward and can be found in A. If $B^{2}=1\!\!1$, such as $B=\text{SWAP}$, the Yang-Baxterization gives a trivial $R(x)$ matrix where the spectral parameter $x$ is only shown in the global phase.
4.2 Yang-Baxter gates from Yang-Baxterization
@@
It follows that they have the same eigenvalues, denoted as
@@
$R_{l,3}(x)=$ $\alpha(x,\lambda_{l,2})B_{l}+\gamma(x,\lambda_{l,1})B^{\dagger}_{l},$ (35b)
@@
Specifically, $R_{I,1}$ is the $R$ matrix of the XXZ model [78]. The braid gate parameter in $R_{I,1}$ corresponds to the anisotropic coupling of the spin chain. The $R_{III,1}$ matrix can also be obtained from the trigonometric limits of elliptic $R$ matrix [79]. Moreover, the $R_{III,1}$ matrix is locally equivalent to the $R$ matrix of the sine-Gordon model [80]. The matrix expressions of $R_{l,1}$ $R_{l,2}$, and $R_{l,3}$, namely total nine $R$ gates, can be found in B.
@@
The optimal gate and pulse constructions of $R_{IV}(\chi)$ with $\varphi_{1}=0$ have been studied in detail in [67].
4.3 Geometric representations of Yang-Baxter gates
@@
The Yang-Baxter gates $R_{I,2}$ and $R_{I,3}$ are SWAP-like gates (exchanging states $|01\rangle$ and $|10\rangle$ with some phases). See the B. Therefore, they have similar geometric representations as $B_{I}$. Specifically, they have the nonlocal parameters
@@
A simple comparison with the nonlocal parameters of $R_{III,1}$ and $R_{I,1}$ in Eq. (38) suggests that $R_{I,1}$ and $R_{III,1}$ have the correspondence through $2\varphi_{1}\leftrightarrow\varphi$ and $2\mu\leftrightarrow\mu$. Therefore, $R_{III,1}$ is also on the faces $OA_{1}A_{2}$ and $OA_{2}A_{3}$ of the two-qubit tetrahedron. See Fig. 4.
@@
4.4 Gate decompositions of Yang-Baxter gates
@@
Recall that the general two-qubit gate can be decomposed into three CNOTs if its three nonlocal parameters are known. Based on the nonlocal parameters of the Yang-Baxter gates presented in the last subsection, we find the optimal gate decomposition (with the minimal number of CNOTs) of $R_{I}$ and $R_{III}$ in terms of the universal gate set $\{R_{z}(\theta),H,\text{CNOT}\}$. See Figs. 5 and 6. For $R_{I}$, we define the parameter $\omega$ as
@@
which only appears in the single-qubit gates of the decomposition. It is also called the $q$ deformation parameter of the $R$ matrix [78]. In practice, if the CNOT is not the native two-qubit gate of the quantum computer, the Yang-Baxter gate which is closest to such native two-qubit gate on the two-qubit tetrahedron has some advantages in realizations. For example, if the quantum computer has the native iSWAP gate, then one can construct the Yang-Baxter gate (with a specific spectral parameter) with only one iSWAP gate with some additional single-qubit gates.
@@
The analysis on $R_{II}$ to be the Clifford gate is the same as $R_{I}$ due to the relation (50). For the case $R_{III}$, we see that there are $T$ gates in the decompositions. See Fig. 6. Therefore, the necessary and sufficient condition for $R_{III}$ to be the Clifford gate becomes $\varphi_{2}\in\{k\pi+\pi/2,k\in\mathbb{Z}\}$, and $a_{III,1},c_{III,1},\varphi_{III,2},\varphi_{III,3}\in\{k\pi/2,k\in\mathbb{Z}\}$. Specifically, the condition $a_{III,1},c_{III,1}\in\{k\pi/2,k\in\mathbb{Z}\}$ (for $R_{III,1}$ to be the Clifford gate) is satisfied if $\mu=0$ or $\varphi_{1}\in\{k\pi/2,k\in\mathbb{Z}\}$. It is similar with $R_{I,1}$. When $\mu=0$ or $\varphi_{1}\in\{k\pi+\pi/2,k\in\mathbb{Z}\}$, we have the Clifford $R_{III,2}$, which is locally equivalent to iSWAP gate; when $\varphi_{1}\in\{k\pi,k\in\mathbb{Z}\}$, we have the Clifford $R_{III,2}$, which is locally equivalent to SWAP gate. Similar results are applied to $R_{III,3}$. The Yang-Baxter gate $R_{IV}$ is Clifford gate if $\varphi_{1}\in\{k\pi,k\in\mathbb{Z}\}$ and $\chi\in\{k\pi/4,k\in\mathbb{Z}\}$, which is locally equivalent to CNOT. We summarize the results in Table 2.
@@ 4.5 Entangling powers
The entangling power of the two-qubit gate defined in Eq. (10) quantifies the ability to generate entanglement from the product state. It can be viewed as the interacting strength between qubits in the brick-wall circuits [54]. Given the nonlocal parameters of $R_{I}$, we numerically calculate its entangling powers in terms of the braid gate parameter $\varphi$ and the spectral parameter $\mu$. See Fig. 7. Specifically, when $\mu=0$ or $\varphi\in\{k\pi,k\in\mathbb{Z}\}$, $R_{I,1,2,3}$ has the zero entangling power, corresponding to the trivial two-qubit gate or SWAP gate. When $\mu\rightarrow\infty$ and $\varphi\in\{k\pi+\pi/2,k\in\mathbb{Z}\}$, $R_{I,1,2,3}$ has the maximal entangling power, corresponding to the iSWAP gate. Note that $\mu\leftrightarrow-\mu$ does not change the magnitude of the nonlocal parameters and the entangling power, therefore we only draw $\mu\geq 0$ in Fig. 7.
@@
Both the phase parameter $\varphi$ and the spectral parameter $\eta$ determine the entangling power of the Yang-Baxter gates. Recall that the spectral parameter is related to the Trotter step of the brick-wall circuit. If we aim to construct states with high entanglement, then the Trotter step should be large with the fixed phase parameter $\varphi=\pi/2$. If we aim to simulate the dynamics of the integrable system with high fidelity, then the Trotter step should be small while avoiding the phase parameter giving large entangling power. This is because the entangled state is more fragile to the noises than the product state.
5 Conclusions
In this work, we have studied the geometric representations of braid and Yang-Baxter gates, where the Yang-Baxter gates are obtained from the Yang-Baxterization of braid gates. Braid and Yang-Baxter gates on the same point of the two-qubit tetrahedron can be transformed from each other through single-qubit gates. In the two-qubit tetrahedron representation, we find that the Yang-Baxter gates can only exist on the edges $A_{2}A_{3}$ and $OA_{1}$, and on the faces $OA_{2}A_{3}$ and $A_{1}A_{2}A_{3}$. Based on their geometric representations, we give their gate decompositions in terms of the universal gate set $\{R_{z}(\theta),H,\text{CNOT}\}$ with the minimal number of CNOT. The braid and Yang-Baxter gates have the maximal entangling power if they are locally equivalent to iSWAP or CNOT. We also find that the Yang-Baxter gates can be the Clifford gate, the matchgate, and the dual unitary gate for specific parameters. It would be interesting to explore the dynamics of the brick-wall circuits composed of such special Yang-Baxter gates. For example, the Yang-Baxter matchgate corresponds to the free-fermion integrable model; the dual unitary Yang-Baxter gate may give special constraints on the speed of information spreading in the integrable circuits.
@@ 5 Conclusions
Data availability statement
All data needed to evaluate the conclusions in the paper are present in the paper. All data that support the findings of this study are included within the article (and any supplementary files).
@@ Appendix A Proof of Theorem 2
$\small R_{I,1}(\mu)\cong\begin{pmatrix}1&0&0&0\\ 0&\dfrac{\sin\varphi}{\sin(\varphi-i\mu)}&\dfrac{-ie^{i\omega}\sinh\mu}{\sin(\varphi-i\mu)}&0\\ 0&\dfrac{-ie^{-i\omega}\sinh\mu}{\sin(\varphi-i\mu)}&\dfrac{\sin\varphi}{\sin(\varphi-i\mu)}&0\\ 0&0&0&1\\ \end{pmatrix},$ (60a)
$\small R_{I,2}(\mu)\cong\frac{1}{\sqrt{\Delta_{I,2}}}\begin{pmatrix}\sinh\left(\frac{1}{2}(\mu+i\varphi)\right)&0&0&0\\ 0&0&e^{i\omega}\sinh\left(\frac{1}{2}(\mu-i\varphi)\right)&0\\ 0&e^{-i\omega}\sinh\left(\frac{1}{2}(\mu-i\varphi)\right)&0&0\\ 0&0&0&\sinh\left(\frac{1}{2}(\mu+i\varphi)\right)\\ \end{pmatrix},$ (60b)
$\small R_{I,3}(\mu)\cong\frac{1}{\sqrt{\Delta_{I,3}}}\begin{pmatrix}\cosh\left(\frac{1}{2}(\mu+i\varphi)\right)&0&0&0\\ 0&0&e^{i\omega}\cosh\left(\frac{1}{2}(\mu-i\varphi)\right)&0\\ 0&e^{-i\omega}\cosh\left(\frac{1}{2}(\mu-i\varphi)\right)&0&0\\ 0&0&0&\cosh\left(\frac{1}{2}(\mu+i\varphi)\right)\\ \end{pmatrix},$ (60c)
$\small R_{II,1}(\mu)\cong\begin{pmatrix}\dfrac{\sin\varphi}{\sin(\varphi-i\mu)}&0&0&\dfrac{-ie^{i\omega}\sinh\mu}{\sin(\varphi-i\mu)}\\ 0&1&0&0\\ 0&0&1&0\\ \dfrac{-ie^{-i\omega}\sinh\mu}{\sin(\varphi-i\mu)}&0&0&\dfrac{\sin\varphi}{\sin(\varphi-i\mu)}\\ \end{pmatrix},$ (60d)
$\small R_{II,2}(\mu)\cong\frac{1}{\sqrt{\Delta_{II,2}}}\begin{pmatrix}0&0&0&e^{i\omega}\sinh\left(\frac{1}{2}(\mu-i\varphi)\right)\\ 0&\sinh\left(\frac{1}{2}(\mu+i\varphi)\right)&0&0\\ 0&0&\sinh\left(\frac{1}{2}(\mu+i\varphi)\right)&0\\ e^{-i\omega}\sinh\left(\frac{1}{2}(\mu-i\varphi)\right)&0&0&0\\ \end{pmatrix},$ (60e)
$\small R_{II,3}(\mu)\cong\frac{1}{\sqrt{\Delta_{II,3}}}\begin{pmatrix}0&0&0&e^{i\omega}\cosh\left(\frac{1}{2}(\mu-i\varphi)\right)\\ 0&\cosh\left(\frac{1}{2}(\mu+i\varphi)\right)&0&0\\ 0&0&\cosh\left(\frac{1}{2}(\mu+i\varphi)\right)&0\\ e^{-i\omega}\cosh\left(\frac{1}{2}(\mu-i\varphi)\right)&0&0&0\\ \end{pmatrix},$ (60f)
@@
$\small R_{III,1}(\mu)\cong\begin{pmatrix}\dfrac{\cosh\mu\cos\varphi_{1}}{\cosh(\mu+i\varphi_{1})}&0&0&-e^{i\varphi_{2}}\dfrac{\sinh\mu\sin\varphi_{1}}{\cosh(\mu+i\varphi_{1})}\\ 0&i\dfrac{\cosh\mu\sin\varphi_{1}}{\sinh(\mu+i\varphi_{1})}&-\dfrac{\sinh\mu\cos\varphi_{1}}{\sinh(\mu+i\varphi_{1})}&0\\ 0&-\dfrac{\sinh\mu\cos\varphi_{1}}{\sinh(\mu+i\varphi_{1})}&i\dfrac{\cosh\mu\sin\varphi_{1}}{\sinh(\mu+i\varphi_{1})}&0\\ e^{-i\varphi_{2}}\dfrac{\sinh\mu\sin\varphi_{1}}{\cosh(\mu+i\varphi_{1})}&0&0&\dfrac{\cosh\mu\cos\varphi_{1}}{\cosh(\mu+i\varphi_{1})}\end{pmatrix},$ (63a)
$\small R_{III,2}(\mu)\cong\frac{1}{\sqrt{\Delta_{III,2}}}\begin{pmatrix}\sinh\mu\cos\varphi_{1}&0&0&e^{i\varphi_{2}}\cosh\mu\sin\varphi_{1}\\ 0&i\cosh\mu\sin\varphi_{1}&-\sinh\mu\cos\varphi_{1}&0\\ 0&-\sinh\mu\cos\varphi_{1}&i\cosh\mu\sin\varphi_{1}&0\\ -e^{-i\varphi_{2}}\cosh\mu\sin\varphi_{1}&0&0&\sinh\mu\cos\varphi_{1}\\ \end{pmatrix},$ (63b)
$\small R_{III,3}(\mu)\cong\frac{1}{\sqrt{\Delta_{III,3}}}\begin{pmatrix}\cosh\mu\cos\varphi_{1}&0&0&-e^{i\varphi_{2}}\sinh\mu\sin\varphi_{1}\\ 0&i\sinh\mu\sin\varphi_{1}&-\cosh\mu\cos\varphi_{1}&0\\ 0&-\cosh\mu\cos\varphi_{1}&i\sinh\mu\sin\varphi_{1}&0\\ e^{-i\varphi_{2}}\sinh\mu\sin\varphi_{1}&0&0&\cosh\mu\cos\varphi_{1}\\ \end{pmatrix},$ (63c)